OpenStudy (anonymous):

Solve √2x+1=x+1 for x. A. No solution B. 0 C. ±2 D. ±3

OpenStudy (anonymous):

You don't wanna try it urself?

OpenStudy (anonymous):

I honestly dont understand how to do this one bit, I do online school, and have to teach myself, so its really difficult

OpenStudy (anonymous):

in the sqrt does in incluse +1 or no

OpenStudy (anonymous):

2x+1 is in the square root

OpenStudy (anonymous):

√2x+1=x+1 1. Move the 1 to the right = (-1) √2x = x+1-1 2. in order to remove √ u have to ^2 the right side. 2x = x^2 2 = x^2 / 2

OpenStudy (anonymous):

Oh is the 1 in the square root?

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

I'll leave it for Karma Ed

OpenStudy (anonymous):

√(2x+1) = x+1 2x +1 = (x+1)^2 I think by now you can do the rest.

OpenStudy (anonymous):

Okay, thank you :)

OpenStudy (anonymous):

Just type here if you get stuck will try to help even more.

OpenStudy (anonymous):

Thanks!